# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

### v1

@@ Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs, which are common in domains such as citation networks, social platforms, and knowledge graphs. GNNs are not inherently interpretable and thus, many explanation methods have been proposed. However, existing explanation methods often struggle to generate interpretable, fine-grained rationales, especially when node attributes include rich natural language. In this work, we introduce Logic, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions. Logic projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure. This enables the LLM to reason about GNN internal representations and produce natural language explanations along with concise explanation subgraphs. Our experiments across four real-world TAG datasets demonstrate that Logic achieves a favorable trade-off between fidelity and sparsity, while significantly improving human-centric metrics such as insightfulness. Logic sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
@@ 1 Introduction
Graph neural networks (GNNs) have witnessed rapid adoption across a wide range of critical real-world applications, including healthcare Zitnik et al. (2018, 2019), drug design Xiong et al. (2021); Liu et al. (2022); Sun et al. (2020), recommender systems Chen et al. (2022), and fraud detection Rao et al. (2021). The high-stakes nature of these domains demands that the predictions made by GNNs be not only accurate but also trustworthy. One powerful approach to building trust in deep learning models is to provide meaningful explanations that justify their predictions Shneiderman (2020). Such explanations, which may highlight important substructures in the input Luo et al. (2020); Schlichtkrull et al. (2021); Yuan et al. (2021), or offer counterfactual scenarios Lucic et al. (2022a); Tan et al. (2022a); Chen et al. (2022). However, generating explanations for GNNs is inherently challenging due to the combinatorial nature of graph data and the joint influence of node attributes and edge connections. These complexities make GNN explainability a non-trivial problem and have led to the development of a wide array of explanation techniques Kakkad et al. (2023a); Guo et al. (2023).
1.1 Need for LLM-based GNN explainers
@@
There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text attribute graphs (TAGs). Essentially, these methods leverage LLMs in different architectural orders to 1) attain rich representations of the input TAGs, task descriptions, and task-specific few-shot examples for the GNN predictors Liu et al. (2024a); Fang et al. (2024), or 2) reason over the TAGs directly Fatemi et al. (2024) or along with GNN-enhanced soft prompts to make predictions Chen et al. (2024a); Tang et al. (2024). However, there is a lack of focus on using LLMs to explain GNNs predictions and clarify the reasoning behind these black-box models.
@@
Traditional GNN explainers aim to identify influential subgraphs or feature subsets that contribute to predictions Ying et al. (2019a); Luo et al. (2020). While effective for graphs with simple node features, these methods struggle in TAGs, where node information is embedded in natural language. Moreover, often these explanations are in the form of subgraphs that are not human-tangible for domain users. LLMs, on the other hand, excel at reasoning over textual content and can generate coherent, human-interpretable rationales. Thus, their integration into GNNs explanation pipelines can improve the interpretability of the GNN models outputs on TAGs.
1.2 Limitations of Existing LLM-based GNN Explainers
@@
The existing frameworks (please see Sec. 2) that use LLMs to build GNN explanations have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainers’ output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or extensive training/fine-tuning of the GNN explainer. (2) Missed GNN internals: Existing approaches fail to directly leverage the rich internal representations of the GNN, making the explanations generic or unfaithful to the internal working of the GNNs. (3) Bias from suboptimal explainers: Involving a pre-existing GNN explainer can bias the reasoning and judgment of the LLM. This is critical when the GNN explainers are suboptimal and they signal noisy information to LLM, while the LLM can potentially better infer and reason the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
1.3 Our Contributions
@@
In this work, we present Logic (Language mOdel-guided GNN Interpretation with Context), a lightweight, post-hoc explanation framework for text-attributed graphs (TAGs) that seamlessly integrates the representational power of GNNs with the reasoning capabilities of LLMs. Our contributions are as follows:
@@
  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by directly using LLMs as interpreters of GNN behavior. Logic generates fine-grained, natural language explanations without relying on handcrafted templates or perturbation-based saliency masks, reducing bias from external explainers.
@@
  • Soft Prompt Integration: Logic aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector. This projection enables GNN features to be treated as soft prompts in a hybrid prompt that blends structural and textual cues, allowing the LLM to reason over the graph’s latent space.
@@
  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Logic produces interpretable subgraphs based on node-level binary support decisions derived from natural language explanations.
@@
  • Training-Agnostic and Plug-and-Play Deployment: Logic does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining. This makes it adaptable for real-world applications.
@@ 2 Related Work
We describe the related studies on the integrated LLM and GNN explainer methods. The explanation capability of LLMs has been applied to graph problems in task-oriented contexts such as explainable recommendation systems Li et al. (2025); Ma et al. (2024). There has been attempts to guide GNNs using the explanations generated by LLMs, for example, in He et al. (2024a), an LLM initially generates node predictions and textual explanations on the reasoning behind the predictions. These explanations then guide a second LLM to produce new node features, which augment the original text features as input for a GNN predictor.
LLM explanation to enhance explainers. LLMs have been utilized to enhance the explanations of pre-existing GNN explainers, for example Zhang et al. (2024) uses LLMs to evaluate explanations of an explainer, its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias. Similarly, to improve the understandability of explanations, Pan et al. (2024) first trains a pseudo-label generator LLM that takes explanations from a pre-existing GNN explainers and generates pseudo-labels for them. The generator is tuned in an expert iteration procedure with tailored objectives for faithfulness and brevity, and the optimized generator generates pseudo-labels to fine-tune an explainer LLM to build end-to-end models.
LLM explanation to enhance understanding. One advantage of LLMs is that they can generate natural language explanations of GNNs, which are more interpretable for humans. He et al. (2024b) also proposes a method for generating counterfactual graph explanations for GNNs, which uses a counterfactual autoencoder (CA)—composed of a contrastively pretrained text encoder based on LLM-generated text pairs and a graph decoder—to construct counterfactual graph topologies by LLM-generated counterfactual text pairs. To address hallucination, the framework incorporates a dynamic feedback mechanism to refine the generated counterfactual text pairs by prompting the LLM back again first-round generated counterfactuals. Concurrent to our work Cedro and Martens (2025) suggests a narrative (story-like) explanation for a GNN’s predictions. Given an LLM, the dataset description, and a GNN model’s predictions, along with the explanation generated by a GNN explainer in the form of subgraphs and feature importance matrices, they ask the LLM to narrate the relationship between the neighboring nodes, the edges connecting them, and the node’s features but fall short of having the LLM as a method to generate the explanations. Our work aims at addressing these limitations and producing a GNN internal representation-aware LLM explanation. We present additional related work in Appendix A.
@@ 3 Problem formulation
A Text-Attributed Graph (TAG) is defined as a graph
@@
Each node $v\in V$ is thus associated with a text document $T(v)$, and the graph structure $E$ represents semantic, relational, or structural links between nodes. TAGs enable learning over both the textual content and the topology of the graph. In this work, we focus on the node classification task.
Node Classification. Given a dataset $\mathcal{G}$ and a graph $G=(V,E,X)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$.
Our method aims at explaining the node classification predictions of a GNN with local subgraphs, i.e, small subgraphs around the node staged for classification Ying et al. (2019a). More formally:
@@
$\Phi(v,\Psi(v,G))=\Phi(v,G)\;\;\forall G\in\mathcal{G},\forall v\in V\cap\mathcal{V}^{\prime}$ (1)
@@
Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ verifies Equation 1 for every node. However, such explainers are rarely useful in practice, as they do not isolate concise or interpretable subgraphs. We therefore evaluate explanations based on their size and fidelity, as formally defined in Section 5.1.
@@
  • It does not necessarily generate concise explanations. Generally, a good factual explanation subgraph is a sparse one Kakkad et al. (2023a); Yuan et al. (2023) which has only the most important features of the graph, i.e., nodes, edges, and node features. This property makes the graph more comprehensible for users.
  • More importantly, this problem definition does not account for human-level understanding of the explanations. For example, retrieving a subgraph as a common form of explanations Ying et al. (2019a); Luo et al. (2020) needs additional information for the domain-experts to understand the results and make them useful in practice.
Nevertheless, these two desirable criteria are indeed compatible and aligned with each other as well as the objective of explainable AI (XAI), which attempts to make black-box AI models more interpretable for humans. Therefore, our goal is to design a method to enhance the understandability of the explanations while making them coherent and concise.
Figure 1: Illustration of Logic’s framework. First, the projector is trained to align GNN node embeddings with the LLM’s embedding space. Next, hybrid prompts are constructed by interleaving projected embeddings (as soft prompts) with natural language tokens. These prompts are then fed into the LLM to produce natural language explanations, which are converted into explanation subgraphs.
4 Our Proposed Method: Logic
To explain the predictions of a trained GNN using an LLM, our method projects node embeddings from the GNN latent space into the LLM embedding space, allowing the LLM to reason about the GNN’s predictions. Rather than relying solely on textual prompts, our approach injects GNN embeddings as soft prompts interleaved with natural language tokens. This enables the LLM to generate context-aware explanations grounded in both the model’s internal representations and human-interpretable features. We present our method in three parts: first, the training of the projector; second, the construction of the hybrid prompt; and third, the generation of the explanation subgraph. An overview of our approach is shown in Figure 1.
@@ 4.1 Projection from the GNN
In order for the LLM to explain the GNN model, we aim to build a framework where the LLM is able to understand how the GNN sees each node of the computation tree when classifying a particular node. To that effect, we introduce a projector that maps nodes’ GNN embedding into tokens.
Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer. The projector aims at both aligning its input with the GNN embeddings and grounding its output in the natural language context of the nodes, hence considering two different levels of embeddings.
@@
Formally, we define a projector as a function:
$\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$
where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings. We write $\Pi(f_{\Phi}(v))=Z_{v}\in\mathbb{R}^{k\times h}$. The projector $\Pi$ is trained to optimize two losses:
@@
    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\left\langle\bar{Z}_{v},\;LLM(v)\right\rangle\right]$
    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output (normalized), and $LLM(v)\in\mathbb{R}^{k}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space. The inner product corresponds to cosine similarity.
@@
    $p_{vu}^{\Phi}=\frac{\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(u)\right\rangle/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(w)\right\rangle/\tau\right)}$
    $p_{vu}^{\Pi}=\frac{\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{u}\right\rangle\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{w}\right\rangle\right)}$
    and $\tau>0$ is a temperature hyperparameter. Both similarity scores are computed using cosine similarity between normalized vectors.
    The overall training objective becomes:
    $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$
    where $1\geq\beta\geq 0$ controls the balance between the two losses.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs (TAGs), which are common in domains such as citation networks, social platforms, and knowledge graphs. GNNs are not inherently interpretable and thus, many explanation methods have been proposed. However, existing explanation methods often struggle to generate interpretable, fine-grained rationales, especially when node attributes include rich natural language. In this work, we introduce Gspell, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions. Gspell projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure. This enables the LLM to reason about GNN internal representations and to produce natural-language explanations, along with concise explanation subgraphs. Our experiments across real-world TAG datasets demonstrate that Gspell achieves a favorable trade-off between fidelity and sparsity, while improving human-centric metrics such as insightfulness. Gspell sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
@@ 1 Introduction
Graph neural networks (GNNs) have witnessed rapid adoption across a wide range of critical real-world applications, including healthcare Zitnik et al. (2018, 2019), drug design Xiong et al. (2021); Liu et al. (2022); Sun et al. (2020), recommender systems Chen et al. (2022), and fraud detection Rao et al. (2021). The high-stakes nature of these domains demands that the predictions made by GNNs be not only accurate but also trustworthy. One powerful approach to building trust in deep learning models is to provide meaningful explanations that justify their predictions Shneiderman (2020). Such explanations, which may highlight important substructures in the input Luo et al. (2020); Schlichtkrull et al. (2021); Yuan et al. (2021); Armgaan et al. (2024), or offer counterfactual scenarios Lucic et al. (2022a); Tan et al. (2022a); Verma et al. (2024); Huang et al. (2023); Fournier and Medya (2025). However, generating explanations for GNNs is inherently challenging due to the combinatorial nature of graph data and the joint influence of node attributes and edge connections. These complexities make GNN explainability a non-trivial problem and have led to the development of a wide array of explanation techniques Kakkad et al. (2023); Guo et al. (2023).
@@
Need for LLM-based GNN explainers. There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text-attributed graphs (TAGs). Essentially, these methods leverage LLMs in different architectural orders to 1) attain rich representations of the input TAGs, task descriptions, and task-specific few-shot examples for the GNN predictors Liu et al. (2024a); Fang et al. (2024), or 2) reason over the TAGs directly Fatemi et al. (2024) or along with GNN-enhanced soft prompts to make predictions Chen et al. (2024a); Tang et al. (2024). However, there is a lack of focus on using LLMs to explain GNN predictions and clarify the reasoning behind these black-box models.
@@
Traditional GNN explainers aim to identify influential subgraphs or feature subsets that contribute to predictions Ying et al. (2019a); Luo et al. (2020). While effective for graphs with simple node features, these methods perform poorly on TAGs, where node information is expressed in natural language. Moreover, these explanations are often presented as subgraphs that are not human-interpretable to users. LLMs, on the other hand, excel at reasoning over textual content and can generate coherent, human-interpretable rationales. Thus, their integration into GNN explanation pipelines can improve the interpretability of the GNN outputs on TAGs.
Limitations of Existing LLM-based GNN Explainers.
@@
The existing frameworks (detailed in Sec. 2) employing LLMs for explaining GNN have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainer’s output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or training/fine-tuning of the GNN explainer. (2) Missed GNN internals: Existing approaches fail to directly leverage the rich internal representations of the GNN, making the explanations generic or unfaithful to the internal working of the GNNs. (3) Bias from suboptimal explainers: Invoking an external GNN explainer can bias the reasoning and judgment of the LLM. This is critical when the GNN explainers are suboptimal, and they signal noisy information to LLM, while the LLM can potentially better infer the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
Our Contributions.
@@
In this work, we present Gspell (GNN Soft Prompted Explanation with LLMs), a lightweight, post-hoc explanation framework for TAGs that integrates the representational power of GNNs with the reasoning capability of LLMs. Our contributions are as follows:
@@
  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by using LLMs as direct interpreters of model behavior. Gspell generates fine-grained, natural language explanations (NLEs) unreliant on handcrafted templates or perturbation-based saliency masks, reducing external explainers’ bias.
@@
  • Soft Prompt Integration: Gspell aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector. This projection enables GNN features to be treated as soft prompts in a hybrid prompt that blends structural and textual cues, allowing the LLM to reason over the graph’s latent space.
@@
  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Gspell produces interpretable NLE subgraphs of node-level support decisions.
@@
  • Training-Agnostic and Plug-and-Play Deployment: Gspell does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining. This makes it adaptable for real-world applications.
@@ 2 Related Work
We describe the related studies on the integrated LLM and GNN explainer methods. The explanation capability of LLMs has been applied to graph problems in task-oriented contexts Li et al. (2025); Ma et al. (2024). There have been attempts to guide GNNs using explanations generated by LLMs; for example, in He et al. (2024a), an LLM generates node predictions and textual explanations of the reasoning behind them.
LLM explanation to enhance explainers. LLMs have been utilized to enhance the explanations of another GNN explainers, for example Zhang et al. (2024) uses LLMs to evaluate explanations of an explainer and its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias. Similarly, to improve the understandability of explanations, Pan et al. (2024) first trains a pseudo-label generator LLM that takes explanations from an external GNN explainers and generates pseudo-labels for them. The generator is tuned in an expert iteration procedure with tailored objectives for faithfulness and brevity, and the optimized generator generates pseudo-labels to fine-tune an explainer LLM to build end-to-end models.
LLM explanation to enhance understanding. One advantage of LLMs is that they can generate NLE for GNNs, which are more human-interpretable. He et al. (2024b) proposes a method for generating counterfactual explanations, which uses an autoencoder to construct counterfactual graph topologies from LLM-generated counterfactual text pairs. To mitigate hallucinations, it employs a dynamic feedback mechanism that prompts the LLM to refine its initial outputs. Lately, Cedro and Martens (2025) proposed prompting an LLM to narrate the explanation subgraphs and the feature importance matrices generated by a GNN explainer, which describe the relationships between neighboring nodes. Our work produces representation-aware LLM explanations without an external explainer. We present additional related work in Appendix A.
@@ 3 Problem formulation
A TAG is defined as a graph
@@
Each node $v\in V$ is thus associated with a text document $T(v)$, and the graph structure $E$ represents semantic, relational, or structural links between nodes. TAGs enable learning from both the textual content and the graph’s topology. In this work, we focus on the node classification task.
Node Classification. Given a dataset $\mathcal{G}$ and a graph $G=(V,E,T)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$, and $x_{v}=f_{e}(T(v))$, where $f_{e}$ is a text encoder.
Our method aims to explain the node classification predictions of a GNN using local subgraphs, i.e., small subgraphs around the node being classified Ying et al. (2019a).
@@
$\Phi(x_{v},\Psi(v,G))=\Phi(x_{v},G)\;\forall G\in\mathcal{G},v\in V\cap\mathcal{V}^{\prime}$
@@
Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ (identity function) verifies Equation 1 for every node. However, such explainers are rarely useful in practice because they do not produce concise, interpretable subgraphs. Therefore, the explanation subgraph size is an important metric, as formally defined in Section 5.1.
@@
  • It does not necessarily generate concise explanations. Generally, a good factual explanation subgraph is sparse Kakkad et al. (2023); Yuan et al. (2023), containing only the most important graph elements: nodes, edges, and node features.
  • More importantly, this problem definition does not account for human-level interpretability. For example, retrieving an explanation subgraph is common Ying et al. (2019a); Luo et al. (2020), but it lacks sufficient information for understanding, which limits its usefulness in practice.
These criteria align with the goals of Explainable AI (XAI). Consequently, we aim to develop a method that enhances interpretability while maintaining conciseness.
Figure 1: Illustration of Gspell’s framework. First, the projector is trained to align GNN node embeddings with the LLM’s embedding space. Next, hybrid prompts are constructed by interleaving projected embeddings (as soft prompts) with natural language tokens. These prompts are then fed to the LLM to produce natural language explanations, which are converted into explanation subgraphs.
4 Our Proposed Method: Gspell
To explain a trained GNN’s predictions, our method projects node embeddings from the GNN’s latent space into the LLM’s embedding space. By injecting these embeddings as soft prompts interleaved with natural language, we enable the LLM to generate context-aware explanations grounded in the model’s internal representations. Our approach consists of three parts: training the projector, constructing the hybrid prompt, and generating the explanation subgraph (Figure 1).
@@ 4.1 Projection from the GNN
To help the LLM interpret the GNN, we developed a framework that aligns the LLM with the GNN’s view of the computation tree. Specifically, we use a projector to map GNN node embeddings directly into the LLM’s token space.
Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E,T)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer. The projector aligns GNN embeddings with natural language context, effectively bridging two distinct embedding spaces.
@@
Formally, we define a projector as a function $\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$, where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings. We write $\Pi(f_{\Phi}(v))=Z_{v}\in\mathbb{R}^{k\times h}$. $\Pi$ optimizes two losses:
@@
    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\text{cos}\left(\bar{Z}_{v},\;LLM(v)\right)\right]$
    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output, $\text{cos}(.,.)$ denotes cosine similarity, and $LLM(v)\in\mathbb{R}^{h}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space.
@@
    $p_{vu}^{\Phi}=\frac{\exp\left(\text{cos}\left(f_{\Phi}(v),f_{\Phi}(u)\right)/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\text{cos}\left(f_{\Phi}(v),f_{\Phi}(w)\right)/\tau\right)}$
    $p_{vu}^{\Pi}=\frac{\exp\left(\text{cos}\left(\bar{Z}_{v},\bar{Z}_{u}\right)\right)}{\sum_{w\in\mathcal{V}}\exp\left(\text{cos}\left(\bar{Z}_{v},\bar{Z}_{w}\right)\right)}$
    and $\tau>0$ is a temperature hyperparameter.
    The overall training objective becomes: $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$ where $1\geq\beta\geq 0$.